OpenStudy (anonymous):

A plane has 90 gallons of fuel. The plane averges 2 2/3 miles per gallon how far can it go?

OpenStudy (anonymous):

240

OpenStudy (anonymous):

ok. multipy?

OpenStudy (anonymous):

Yeah/

OpenStudy (anonymous):

2 2/3 = 8/3 distance travel : 90(gallons) x 8/3 ( mile/gallon) = 240 mile